### v1

# compat=1.1.0

Abstract
We present a dequantization algorithm for the Vlasov–Poisson (VP) system, termed the dequantized particle algorithm, by systematically dequantizing the underlying many-body quantum theory. Starting from the second-quantized Hamiltonian description, we derive a finite-dimensional dequantized system and show that it furnishes a structure-preserving discretization of the Schrödinger–Poisson (SP) equations. Through the Wigner or Husimi transformations, this discretization provides an efficient approximation of the VP system when quantum effects are negligible. Unlike conventional structure-preserving algorithms formulated in 6D phase space, this dequantized particle algorithm operates in 3D configuration space, potentially offering more compact and efficient representations of physical information under appropriate conditions. A numerical example of the classical nonlinear two-stream instability, simulated using merely 97 dequantized particles, demonstrates the efficiency, accuracy, and conservation properties of the algorithm and confirms its potential as a foundation for developing quantum and quantum-inspired classical algorithms for kinetic plasma dynamics.
The dynamics of a collection of $N$ identical bosonic charged particles interacting through Coulomb potential can be described by the Hamiltonian
$H=\sum_{i=1}^{N}\frac{\boldsymbol{p}_{i}^{2}}{2m}+\frac{1}{2}\sum_{i,j=1}^{N}\frac{q^{2}}{\left|\boldsymbol{x}_{i}-\boldsymbol{x}_{j}\right|}\,,$ (1)
where $q$ and $m$ are the charge and mass of the particles. In quantum many-body theory, $\boldsymbol{p}_{i}=i\hbar\partial/\partial\boldsymbol{x}_{i}$ and the system is governed by the Schrödinger equation
$\frac{\partial}{\partial t}\psi(\boldsymbol{x}_{1},...,\boldsymbol{x}_{N},t)=\frac{1}{i\hbar}H\psi(\boldsymbol{x}_{1},...,\boldsymbol{x}_{N},t)\,.$ (2)
In classical theory, $\boldsymbol{p}_{i}$ $(i=1,...,N)$ are independent variables of the $N$-particle phase space $\boldsymbol{z}_{i}=(\boldsymbol{x}_{i},\boldsymbol{p}_{i})$ $(i=1,...,N)$, and the dynamics are governed by Hamilton’s equation
$\frac{d\boldsymbol{z}_{i}}{dt}=\left\{\boldsymbol{z}_{i},H\right\}.$ (3)
The many-body quantum theory recovers the classical theory when the characteristic kinetic energy of the particles is much larger than the characteristic Coulomb potential, $\boldsymbol{p}_{i}^{2}/2m\gg q^{2}/\left|\boldsymbol{x}_{i}-\boldsymbol{x}_{j}\right|$. To simplify notation, we will overload the symbol $H$ to denote both quantum operators and classical Hamiltonian functions, depending on the context.
In classical theory, when the correlation between particles is negligible, the system can be described by a one-particle distribution function $f(\boldsymbol{x},\boldsymbol{v},t)$, which satisfies the Vlasov-Poisson (VP) equations,
$\frac{\partial f}{\partial t}$ $+\boldsymbol{v}\cdot\frac{\partial f}{\partial\boldsymbol{x}}-\frac{q}{m}\boldsymbol{\nabla}\phi\cdot\frac{\partial f}{\partial\boldsymbol{v}}=0\,,$ (4)
$\nabla^{2}\phi$ $=-4\pi q\left(\int d^{3}vf-\frac{N}{L^{3}}\right)\,,$ (5)
where $\boldsymbol{v}\equiv\boldsymbol{p}/m$, $L$ is the linear size of the system, and a uniform background charge density is assumed such that the system is overall charge neutral.
The VP system has many important applications in plasma physics, fusion energy, and stellar dynamics. Numerical algorithms for the VP equations—and the more general Vlasov–Maxwell (VM) equations—have long been an active area of research in these fields. In particular, recent years have witnessed significant progress in the development of classical structure-preserving algorithms for the VP and VM systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], quantum algorithms for the linear VP and VM systems [22, 23, 24, 25, 26] and the nonlinear VP system [27], and quantum-classical hybrid algorithm for the nonlinear VM system [28]. However, quantum algorithms currently available for these systems are not yet practical on present-day quantum hardware, especially when compared to state-of-the-art structure-preserving algorithms running on classical computers. Nevertheless, it is anticipated that quantum simulations will eventually demonstrate advantages for certain classes of problems on future quantum computers. Moreover, studying quantum algorithms may also lead to efficient quantum-inspired classical algorithms, which are referred to as dequantization algorithms in this context.
In the present study, we develop such a dequantization algorithm for the nonlinear VP system, termed the dequantized particle algorithm, by dequantizing the many-body quantum theory specified by Eq. (2). This procedure yields a finite-dimensional, structure-preserving discretization of the Schrödinger–Poisson (SP) equations, whose solutions approximate those of the VP system when quantum effects are negligible. Compared with established structure-preserving algorithms for the VP system in 6D phase space, the dequantized particle algorithm offers an alternative formulation in 3D configuration space. Moreover, it provides a two-way bridge that can also be used to develop quantum algorithms for the VP system.
We start from the quantum many-body system governed by Eq. (2). The Hamiltonian of the system can be equivalently cast into the following form using the technique of second quantization [29, 30],
$H$ $=\sum_{ij}H_{ij}\hat{a}_{i}^{\dagger}\hat{a}_{j}+\frac{1}{2}\sum_{ijln}\hat{a}_{i}^{\dagger}\hat{a}_{j}^{\dagger}V_{ijln}\hat{a}_{l}\hat{a}_{n},$ (6)
$H_{ij}$ $\equiv-\int\psi_{i}^{*}(\boldsymbol{x})\frac{\hbar^{2}}{2m}\nabla^{2}\psi_{j}(\boldsymbol{x})d^{3}\boldsymbol{x},$ (7)
$V_{ijln}$ $\equiv\int\psi_{i}^{*}(\boldsymbol{x}_{1})\psi_{j}(\boldsymbol{x}_{1})\frac{q^{2}}{\left|\boldsymbol{x}_{1}-\boldsymbol{x}_{2}\right|}\psi_{l}^{*}(\boldsymbol{x}_{2})\psi_{n}(\boldsymbol{x}_{2})\,d^{3}\boldsymbol{x}_{1}d^{3}\boldsymbol{x}_{2}.$ (8)
Here, $\psi_{n}(\boldsymbol{x})\,\,(n\in\mathbb{N})$ is a set of suitable basis functions on one-particle configuration space, and $\hat{a}_{n}^{\dagger}$ and $\hat{a}_{n}\,\,(n\in\mathbb{N})$ are the corresponding creation and annihilation operators. Since the free particle Hamiltonian is $\hbar^{2}\nabla^{2}/2m$, we will take $\psi_{n}(\boldsymbol{x})=\exp(-i\boldsymbol{k}_{n}\cdot\boldsymbol{x})L^{-3/2}$ with $\boldsymbol{k}_{n}=2\pi n/L$. Note that $n$ is a vector index. For this choice of basis functions,
$H$ $=\sum_{l}\frac{k_{l}^{2}\hbar^{2}}{2m}\hat{a}_{l}^{{\dagger}}\hat{a}_{l}+\sum_{l,n,g\neq 0}\frac{4\pi q^{2}}{L^{3}k_{g}^{2}}\hat{a}_{l-g}^{{\dagger}}\hat{a}_{n+g}^{{\dagger}}\hat{a}_{l}\hat{a}_{n}\thinspace,$ (9)
where use has been made of
$V(\boldsymbol{k})\equiv\int\frac{q^{2}}{\left|\boldsymbol{x}\right|}\exp(-i\boldsymbol{k}\cdot\boldsymbol{x})d^{3}\boldsymbol{x}=\frac{4\pi q^{2}}{k^{2}}.$ (10)
The dynamics of the system is governed by
$\dot{\hat{a}}_{j}=\frac{1}{i\hbar}[\hat{a}_{j},H]=\frac{-i\hbar k_{j}^{2}}{2m}\hat{a}_{j}+\frac{2\pi q^{2}}{i\hbar L^{3}}\sum_{l,g\neq 0}\frac{1}{k_{g}^{2}}\hat{a}_{l}^{\dagger}\hat{a}_{l-g}\hat{a}_{j+g}\thinspace.$ (11)
The first term on the right hand side of Eq. (9) is the free-particle operator, and the second term represents the interaction between two particles through Coulomb interaction, as illustrated in Fig. 1.
[…]
Figure 1: Two interacting particles exchange momentum via a photon that carries momentum $\hbar\boldsymbol{k}_{g}$.
For practical calculations, it is necessary to truncate the system for a finite number of $\hat{a}_{j}$. However, it is essential that the truncated system preserves all the important structures of the original system, including Hermiticity and the conservation of particles, energy, and momentum. To this end, we select a finite set $J$ of vector indices and discretize (truncate) the system as follows,
$H_{d}$ $\equiv\sum_{l\in J}\frac{k_{l}^{2}\hbar^{2}}{2m}\hat{a}_{l}^{{\dagger}}\hat{a}_{l}+\sum_{l,n,l-g,n+g\in J}^{g\neq 0}\frac{4\pi q^{2}}{L^{3}k_{g}^{2}}\hat{a}_{l-g}^{{\dagger}}\hat{a}_{n+g}^{{\dagger}}\hat{a}_{l}\hat{a}_{n}\thinspace,$ (12)
$\dot{\hat{a}}_{j}=\frac{1}{i\hbar}[\hat{a}_{j},H_{d}]=\frac{-k_{j}^{2}i\hbar}{2m}\hat{a}_{j}+\frac{2\pi q^{2}}{i\hbar L^{3}}\sum_{l,l-g,j+g\in J}\frac{1}{k_{g}^{2}}\hat{a}_{l}^{\dagger}\hat{a}_{l-g}\hat{a}_{j+g}\thinspace,\,\,\,\,j\in J.$ (13)
It is crucial to note that we require $l,n,l-g,n+g\in J,$ but $g\in J$ is not guaranteed or required. Equations (12) and (13) include the term $k_{g}^{2}$, for which the corresponding $\hat{a}_{g}$ may not exist in the discretized system $H_{d}.$ Physically, $\hbar\boldsymbol{k}_{g}$ is momentum of the gauge photon mediating the interaction between two particles, see Fig. 1.
It is straightforward to see that because of this specific truncation method, $H_{d}$ remains Hermitian after the truncation. Obviously, $H_{d}$ is conserved for the discretized system, i.e., $\left[H_{d},H_{d}\right]=0.$ From a physics perspective, the discrete system also conserves the total number of particles, $N_{d}\equiv\sum_{l\in J}\hat{a}_{l}^{{\dagger}}\hat{a}_{l}$, because in the expression of $H_{d}$, the number of creation operators exactly balances the number of annihilation operators. Mathematically, we have $\left[H_{d},N_{d}\right]=0,$ which can be proved by applying the Leibniz rule as follows. For every quartet $l,m,l-g,n+g\in J$,
$\left[\hat{a}_{l-g}^{{\dagger}}\hat{a}_{n+g}^{{\dagger}}\hat{a}_{l}\hat{a}_{n},\sum_{j\in J}\hat{a}_{j}^{{\dagger}}\hat{a}_{j}\right]$ $=\left[\hat{a}_{l-g}^{{\dagger}},\sum_{j\in J}\hat{a}_{j}^{{\dagger}}\hat{a}_{j}\right]\hat{a}_{n+g}^{{\dagger}}\hat{a}_{l}\hat{a}_{n}+\hat{a}_{l-g}^{{\dagger}}\left[\hat{a}_{n+g}^{{\dagger}},\sum_{j\in J}\hat{a}_{j}^{{\dagger}}\hat{a}_{j}\right]\hat{a}_{l}\hat{a}_{n}$
$+\hat{a}_{l-g}^{{\dagger}}\hat{a}_{n+g}^{{\dagger}}\left[\hat{a}_{l},\sum_{j\in J}\hat{a}_{j}^{{\dagger}}\hat{a}_{j}\right]\hat{a}_{n}+\hat{a}_{l-g}^{{\dagger}}\hat{a}_{n+g}^{{\dagger}}\hat{a}_{l}\left[\hat{a}_{n},\sum_{j\in J}\hat{a}_{j}^{{\dagger}}\hat{a}_{j}\right]$
$=(-1-1+1+1)\hat{a}_{l-g}^{{\dagger}}\hat{a}_{n+g}^{{\dagger}}\hat{a}_{l}\hat{a}_{n}=0\,.$ (14)
Similarly, the momentum operator, $P_{d}\equiv\sum_{n\in J}k_{n}\hat{a}_{n}^{{\dagger}}\hat{a}_{n},$ of the discrete system is conserved because for each term $\hat{a}_{l-g}^{{\dagger}}\hat{a}_{n+g}^{{\dagger}}\hat{a}_{l}\hat{a}_{n}$ in $H_{d}$, $\hat{a}_{l}\hat{a}_{n}$ destroys as much momentum as $\hat{a}_{l-g}^{{\dagger}}\hat{a}_{n+g}^{{\dagger}}$ creates. This can be mathematically expressed as that for each quartet $l,n,l-g,n+g\in J$, akin to the calculation in Eq. (14), we have
$\left[\hat{a}_{l-g}^{{\dagger}}\hat{a}_{n+g}^{{\dagger}}\hat{a}_{l}\hat{a}_{n},\sum_{j\in J}k_{j}\hat{a}_{j}^{{\dagger}}\hat{a}_{j}\right]=\left(-k_{l-g}-k_{n+g}+k_{l}+k_{n}\right)\hat{a}_{l-g}^{{\dagger}}\hat{a}_{n+g}^{{\dagger}}\hat{a}_{l}\hat{a}_{n}=0,$ (15)
which proves $\left[H_{d},P_{d}\right]=0.$
The discretized system $H_{d}$ is now structure-preserving and can, in principle, be simulated on a quantum computer. However, in the present study, travel along the other direction—we dequantize $H_{d}$ into a classical Hamiltonian function,
$H_{d}$ $=H_{0}+H_{1},\,\,$ (16)
$H_{0}$ $=\sum_{l\in J}\frac{k_{l}^{2}\hbar^{2}}{2m}a_{l}^{*}a_{l},\thinspace\thinspace H_{1}=\sum_{l,n,l-g,n+g\in J}^{g\neq 0}\frac{4\pi q^{2}}{L^{3}k_{g}^{2}}a_{l-g}^{*}a_{n+g}^{*}a_{l}a_{n}\thinspace,$ (17)
where $\hat{a}_{l}^{\dagger}$ and $\hat{a}_{l}$ have been dequantized into c-numbers $a_{l}^{*}$ and $a_{l}$, respectively. In the dequantized theory, $H_{d}$ becomes the total energy, which splits into $H_{0}$, the kinetic energy of dequantized particles, and $H_{1},$ the potential energy between the dequantized particles. The equation of motion for $a_{j}\,(j\in J)$ is
$\dot{a}_{j}=\frac{1}{i\hbar}\left\{a_{j},H_{d}\right\}=\frac{-k_{j}^{2}i\hbar}{2m}a_{j}+\frac{2\pi q^{2}}{i\hbar L^{3}}\sum_{l,l-g,j+g\in J}^{g\neq 0}\frac{1}{k_{g}^{2}}a_{l}^{*}a_{l-g}a_{j+g}\thinspace.$ (18)
[…]
What physics does the finite-dimensional dequantized system (16) and (18) describe? It turns out it represents a structure-preserving discretization of the Schrödinger-Poisson (SP) equations,
$\frac{\partial\psi}{\partial t}$ $=\frac{1}{i\hbar}\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+q\phi\right]\psi\,,$ (19)
$\nabla^{2}\phi$ $=-4\pi q\left(\psi^{*}\psi-\frac{N}{L^{3}}\right)\,.$ (20)
Here, $\psi(\boldsymbol{x},t)$ is the wave function of one particle. To establish this fact, we work backward by finding a specific discretization scheme for Eqs. (19) and (20) such that the resulting discretization agrees with Eqs. (16) and (18).
To achieve this goal, we first reformulate Eqs. (19) and (20) as a Hamiltonian system defined by the Hamiltonian functional
$H(\psi,\psi^{*})$ $=\int d^{3}\boldsymbol{x}\left\{\psi^{*}\left[-\frac{\hbar^{2}}{2m}\nabla^{2}+\frac{1}{2}q\phi\right]\psi-\frac{N}{2}\frac{q}{L^{3}}\phi\right\}\,.$ (21)
Here, it is understood that $\phi(\boldsymbol{x},t)$ is a function determined by $\psi(\boldsymbol{x},t)$ via Eq. (20) at every $t,$ since $\phi(\boldsymbol{x},t)$ cannot be cast as an independent dynamical field. It can be shown that Eq. (19) is equivalent to
$\frac{\partial\psi}{\partial t}=\frac{1}{i\hbar}\frac{\delta H}{\delta\psi^{*}}\,.$ (22)
Here, $\delta H/\delta\psi^{*}$ denotes functional derivative. Note that $H$ depends on $\psi$ and $\psi^{*}$ explicitly and also implicitly through $\phi(\boldsymbol{x},t).$ Both contributions need to be included in the calculation of $\delta H/\delta\psi^{*}$.
We discretize the Hamiltonian in Eq. (21) in terms of Fourier components. Let
$\psi(\boldsymbol{x},t)$ $=\sum_{l}a_{l}(t)\exp(ik_{l}\cdot\boldsymbol{x})L^{-3/2}\,,$ (23)
$\phi(\boldsymbol{x},t)$ $=\sum_{l}\phi_{l}(t)\exp(ik_{l}\cdot\boldsymbol{x})L^{-3/2}\,,$ (24)
From Eq. (20), we have
$\phi_{l}=4\pi q\frac{1}{k_{g}^{2}}\sum_{n}a_{n}^{*}a_{g+n},\ (g\neq 0).$ (25)
The 00-th component of $\phi$ is an arbitrary constant, $\phi_{0}=\text{const.}\,.$. . In terms of $a_{l}(t),$ we find
$H(\psi,\psi^{*})=\sum_{l}\frac{k_{l}^{2}\hbar^{2}}{2m}a_{l}^{*}a_{l}+\sum_{l,n,g\neq 0}\frac{4\pi q^{2}}{L^{3}k_{g}^{2}}a_{l-g}^{*}a_{n+g}^{*}a_{l}a_{n}\thinspace,$ (26)
which reduces to $H_{d}$ in Eq. (16) once the truncation condition $l,n,l-g,l+g\in J$ is imposed. Thus, we have proved that the dequantized system (16) and (18) is a finite-dimensional, structure-preserving discretization of the SP system.
The correspondence between VP and SP systems had been established [31, 32, 33, 34, 35, 36, 37, 38]. In its simplest form, the phase space distribution $f(\boldsymbol{x},\boldsymbol{v},t)$ in the VP system can be identified with the Wigner function constructed from the wave function $\psi(\boldsymbol{x},t)$ in the SP system,
$f(\boldsymbol{x},\boldsymbol{v},t)=\frac{1}{(2\pi\hbar/m)^{3}}\int_{\mathbb{R}^{3}}d^{3}\boldsymbol{y}\,\exp\!\Big{(}-\frac{i}{\hbar}m\boldsymbol{v}\cdot\boldsymbol{y}\Big{)}\,\psi^{*}\Big{(}\boldsymbol{x}+\frac{\boldsymbol{y}}{2}\Big{)}\,\psi\Big{(}\boldsymbol{x}-\frac{\boldsymbol{y}}{2}\Big{)}.$ (27)
It is worth emphasizing that the distribution function $f(\boldsymbol{x},\boldsymbol{v},t)$ constructed this way is not semi-positive definite—a trade-off for the simplicity of its form. The negative values of $f(\boldsymbol{x},\boldsymbol{v},t)$ arise from quantum effects [32, 33, 39, 35], which are assumed to be small for the classical problems under study. Husimi [31] showed that the Wigner function can be smoothed with a minimum-uncertainty Gaussian kernel to produce a non-negative phase space function now known as the Husimi function. Cartwright [34] extended this idea by proving that if the Wigner function is smoothed with any kernel whose width is not smaller than that of the minimum-uncertainty state, the resulting distribution remains non-negative for all quantum states. Together, these works establish that by appropriate coarse-graining, the Wigner transformation can yield a non-negative quasi-probability distribution. With this correspondence, Eqs. (16) and (18) furnish a finite-dimensional, structure-preserving discretization of the VP system. This dequantized particle algorithm offers several distinct advantages over conventional classical approaches. Most notably, it evolves an equation system defined in 3D configuration space, whereas classical algorithms operate in 6D phase space. This dimensional reduction suggests that the dequantized particle algorithm can represent the physical system’s information more efficiently under suitable conditions.
Finally, we comment on the computational complexity of the dequantized particle algorithm. For a system with $M$ dequantized particles, the complexity of one-step time-advance map is ostensibly $M^{3},$ judging from summation of the $a_{l}^{*}a_{l-g}a_{j+g}$ terms in Eq. (18). However, this summation consists of convolution summations, which can be efficiently calculated in $M\log M$ operations, akin to the standard spectrum and PIC algorithms.
We now present a numerical example of the dequantized particle algorithm using the test problem of 1D two-stream instability. For numerical calculations, we normalize length by $L,$ time by $\omega_{p}$, and energy by $mL^{2}\omega_{p}^{2}$. Here, $\omega_{p}\equiv\sqrt{4\pi Nq^{2}/L^{3}m}$ is the plasma frequency. Specifically, the normalized variables are
$\bar{\boldsymbol{x}}=\boldsymbol{x}/L,\,\,\,\bar{t}=t\omega_{p},\thinspace\thinspace\thinspace\bar{\psi}=\psi L^{3/2}N^{-1/2},\,\,\,\bar{a}_{j}=a_{j}N^{-1/2},\,\,\,\bar{\phi}=\frac{q\phi}{mL^{2}\omega_{p}^{2}},\,\,\,\bar{H}=\frac{H}{\omega_{p}}.$ (28)
In normalized variables, Eq. (18) is
$\dot{a}_{j}=\frac{\delta}{2i}a_{j}k_{j}^{2}+\frac{1}{2i\delta}\sum_{l,l-g,j+g\in J}^{g\neq 0}\frac{1}{k_{g}^{2}}a_{l}^{*}a_{l-g}a_{j+g}\thinspace,\,\,\,j\in J\,.$ (29)
Here, the over bar “$\bar{\thinspace\thinspace\,}$” for normalized variables has been dropped for easy notation and
$\delta\equiv\frac{\hbar}{L^{2}\omega_{p}m}=\left(\frac{\hbar}{Lmv_{th}}\right)\left(\frac{\lambda_{D}}{L}\right)$ (30)
is a dimensionless parameter measuring the quantum effect using the characteristics of a classical plasma. Here, $v_{th}$ is the thermal velocity of the plasma and $\lambda_{D}$ is the Debye length. For classical dynamics, $\delta\ll 1.$
Figure 2: Nonlinear two-stream instability simulated with 97 dequantized particles. Shown is the distribution function $f(x,v,t)$ at $t=0,16,20,22,23,25$, constructed using Eq. (27) from $a_{j}\,(-48\leq j\leq 48).$ Phase-space vortex structures develop during the instability.
Figure 3: The wave function $\psi(x,t)$ at $t=0,16,20,22,23$, and 25 is constructed from the 97 dequantized particles $a_{j}\,(-48\leq j\leq 48).$ In the linear phase, the wave function is dominated by a small number of dequantized particles, whereas in the nonlinear phase, more dequantized particles participate in the dynamics.
Figure 4: The photon field $\phi(x,t)$ at $t=0,16,20,22,23,$ and $25$ is constructed from the 97 dequantized particles $a_{j}\,(-48\leq j\leq 48).$ In the linear phase, the potential energy $H_{1}$ increases monotonically, while in the nonlinear phase, the mediating photon field deviates significantly from being monochromatic. To illustrate the structure of the field during the instability, different scales are used for the different panels.
To load in an initial two-stream distribution with peaked population at $v=\pm V_{0},$ we follow Ref. [35] to let
$\psi(x,t=0)=\sqrt{2}\cos\left(\frac{xV_{0}}{\delta}\right)\exp\left(-\frac{i\epsilon V_{0}}{\delta k}\cos(kx)\right),$ (31)
where an initial perturbation at wavelength $k$ and amplitude $\epsilon$ has also been included. Displayed in Figs. 1-5 are simulation results for the case of $(V_{0},\delta,k,\epsilon)=(0.04854,1.9\times 10^{-4},4\pi,5\times 10^{-4})$ with 97 dequantized particles, i.e., $J=\left\{n\in\mathbb{N}|-48\leq n\leq 48\right\}.$ The simulation was run from $t=0$ to $30$. The time history of $f(x,v,t)$, constructed using Eq. (27), is displayed in Fig. 2, which illustrates the development of phase-space vortex structures during the instability. As noted previously, the negative values of $f(x,v,t)$ are associated with the quantum effects [32, 33, 39, 35], which manifest as small scale structures with wavelengths proportional to $\delta$ in this problem. If preferred, the non-negative Husimi function [31, 34] can be used as an alternative. By The time history of $\psi(x,t)$ and $\phi(x,t)$ are shown in Figs. 3 and 4, respectively. The two-stream instability grows exponentially in the linear phase and saturates approximately at $t=25.$ The growth rate at the linear phase is measured to be $\gamma=0.3493$ [Fig. 5(a)], which agrees well with the theoretical growth rate $\gamma_{\text{theory}}=0.3536,$ solved for from the analytical dispersion relation
$\frac{1}{2(\omega+kV_{0})^{2}}+\frac{1}{2(\omega-kV_{0})^{2}}=1.$ (32)
The time history of the total energy, number of particles, and total momentum are plotted in Fig. 5. All three quantities are well conserved to high precision. In the linear phase, the wave function is dominated by a small number of dequantized particles, the kinetic energy $H_{0}$ decreases monotonically, the potential energy $H_{1}$ increases monotonically, whereas the total energy $H_{d}$ is conserved. In the nonlinear phase, more dequantized particles participate in the dynamics, the mediating photon field deviates significantly from being monochromatic, and both $H_{0}$ and $H_{1}$ oscillate at approximately the same level.
Figure 5: Nonlinear two-stream instability simulated with 97 dequantized particles. (a) Growth rate; (b) time history of energy; (c) number of particles; (d) momentum. The initial momentum is zero for the two-stream distribution, and the momentum variation is measured by $\sqrt{mH_{d}}$. Energy, number of particles, and momentum are well conserved to high precision.
In conclusion, we have developed a dequantization algorithm for the VP system, termed the dequantized particle algorithm, by dequantizing the many-body quantum Hamiltonian in its second-quantized form. The resulting finite-dimensional system preserves the key conservation laws and furnishes a structure-preserving discretization of the SP equations, which connects naturally to the VP system through the Wigner transformation. This formulation enables an alternative simulation strategy in 3D configuration space rather than the full 6D phase space for the VP system. The example of the two-stream instability using 97 dequantized particles, which correspond to the Fourier modes of the VP system, confirms that the algorithm accurately captures the expected physical behavior and conserves invariants to high precision. These results highlight the potential of the dequantization approach to provide new pathways for developing efficient, structure-preserving classical algorithms as well as quantum algorithms for plasma kinetic simulations [40, 41, 42, 43, 44, 45, 23, 46, 47, 48].
Acknowledgements.
This research is supported by the U.S. Department of Energy (DE-AC02-09CH11466). J. M. Molina acknowledges support from the National Science Foundation Graduate Research Fellowship under Grant No. KB0013612.